sir please help me to solve this question
write a pythagoren triplet whose smallest member is 8.

Found 2 solutions by Alan3354, MathLover1:
Answer by Alan3354(69443)

8, 15, 17
-------------
8^2 = 64
64/2 = 32
32 = 15 + 17

Answer by MathLover1(20849)

it is ( 8, 15, 17)